### v1

# Regge trajectories for the hidden bottom tetraquarks in the diquark picture

Abstract
We apply the newly proposed Regge trajectory relation for the heavy-light diquarks alongside the Regge trajectory formula for the heavy-heavy systems to investigate the hidden bottom tetraquarks $(bq)(\bar{b}\bar{q}^{\prime})$ $(q,\,q^{\prime}=u,\,d,\,s)$. The Regge trajectory formulas for the hidden bottom tetraquarks are obtained. The masses of the $\lambda$-mode excited states and the $\rho$-mode excited states of the hidden bottom tetraquarks are estimated. Additionally, the $\lambda$-trajectories and the $\rho$-trajectories of the tetraquarks with hidden bottom are discussed. We find that the behavior of the $\rho$-trajectories is different from that of the $\lambda$-trajectories. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ $(x_{\rho}=n_{r},\,l)$ because the $\rho$-mode excitations are in diquark and antidiquark, with both the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ being the heavy-light systems. Conversely, the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ $(x_{\lambda}=N_{r},\,L)$ because the $\lambda$-mode excitations occure between diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$, with the the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ forming a heavy-heavy system $(bq)(\bar{b}\bar{q}^{\prime})$. Moreover, not only the $\lambda$-trajectories but also the $\rho$-trajectories for the hidden bottom tetraquarks are concave downwards in the $(M^{2},\,x)$ plane.
Regge trajectory, tetraquark, mass
I Introduction
As a type of exotic hadrons, tetraquarks will enhance the study of hadrons and will provide new probes for the understanding of QCD Jaffe:1976ig ; Jaffe:1976ih ; pdg:2024 ; Gross:2022hyw ; Bicudo:2022cqi ; Liu:2019zoy ; Jaffe:2004ph ; Esposito:2016noz . The hidden bottom tetraquarks, composed of two bottom quarks and two light quarks, have been studied through various methods. These approaches include the relativistic quark model based on the quasipotential approach in QCD Ebert:2008se ; Faustov:2021hjs , the non-relativistic potential model Tiwari:2022azj ; Lundhammar:2020xvw , the spinless Salpeter equation Ferretti:2020ewe ; Giannuzzi:2023vrx , the relativistic four-body Faddeev-Yakubovsky equation Hoffer:2024alv , the dynamical diquark model Giron:2020qpb , the homogeneous Lippmann-Schwinger integral equation Hadizadeh:2015cvx , the effective Hamiltonian in the adopted chromomagnetic interaction model Wu:2018xdi , the chiral quark model Liu:2022vyy , the sum rules Chen:2017dpy ; Wang:2019mxn , the relativistic four-quark equations in the framework of coupled-channel formalism Gerasyuta:2008cb , among others.
The Regge trajectory is one of the effective approaches widely used in the study of hadron spectra Regge:1959mz ; Chew:1961ev ; Chew:1962eu ; Nambu:1978bd ; Collins:1971ff ; Inopin:1999nf ; Irving:1977ea . Ref. Selem:2006nd gives a compact and predictive unified picture of mesons, baryons, and tetraquarks with the loaded flux-tube model. Ref. Nielsen:2018uyn presents relations between the Regge trajectories of mesons, baryons, and tetraquarks based on the Holographic light front QCD. In Ref. Sonnenschein:2018fph , the masses of mesons, baryons, glueballs and tetraquarks are predicted by the Regge trajectories obtained from the Holography Inspired Stringy Hadron model. In Ref. MartinContreras:2020cyg , mesons, hybrid mesons and multiquark states are discussed by the nonlinear Regge trajectories in the model AdS/QCD with dilaton. In Refs. Chen:2023web ; Chen:2023djq , we discuss the universal descriptions of the heavy-heavy and heavy-light systems (mesons, baryons and tetraquarks) by using the Regge trajectories  [^1] A Regge trajectory of hadrons is assumed to be written as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$, where $M$ is mass of the bound state, $l$ is the the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, the plots in the $(M,\,x)$ plane, in the $(M,\,(x-c_{0})^{1/{\nu}})$ plane, in the $(M^{2},\,x)$ plane, in the $((M-m_{R})^{2},\,x)$ plane or in the $((M-m_{R})^{1/{\nu}},\,x)$ plane are all called the Chew-Frautschi plots. The Regge trajectories can be plotted in these different planes. . In the diquark picture, the hidden bottom tetraquarks are composed of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$ ($q,q^{\prime}=u,d,s$). By applying the Regge trajectory relation for the heavy-light diquarks proposed in Ref. Chen:2023cws and the Regge trajectory formula for the heavy-heavy systems Chen:2023djq , we investigate the hidden bottom tetraquarks. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks with the hidden bottom are estimated, and the $\lambda$-trajectories and $\rho$-trajectories are discussed.
The paper is organized as follows: In Sec. II, the Regge trajectory relations are obtained from the spinless Salpeter equation. In Sec. III, we investigate the Regge trajectories for the hidden bottom tetraquarks. The conclusions are presented in Sec. IV.
II Regge trajectory relations for the hidden bottom tetraquarks
In this section, with the help of the diquark Regge trajectories Chen:2023cws ; Feng:2023txx ; Chen:2023ngj , we present the Regge trajectory relations for the hidden bottom tetraquarks, which can discuss both the $\lambda$-trajectories and the $\rho$-trajectories.
II.1 Preliminary
In the diquark picture, tetraquarks consist of one diquark and one antidiquark, see Fig. 1. $\rho_{1}$ ($\rho_{2}$) separates the quarks (antiquarks) in the diquark (antidiquark), and $\lambda$ separates the diquark and the antidiquark. There exist three excited modes: the $\rho_{1}-$mode involves the radial and orbital excitation in the diquark, the $\rho_{2}-$mode involves the radial and orbital excitation in the antidiquark, and the $\lambda-$mode involves the radial or orbital excitation between the diquark and antidiquark. Consequently, there exist three series of Regge trajectories: two series of $\rho$-Regge trajectories and one series of $\lambda$-Regge trajectories.
In $SU_{c}(3)$, there is attraction between quark pairs $(qq^{\prime})$ in the color antitriplet channel and this is just twice weaker than in the color singlet $q\bar{q}^{\prime}$ in the one-gluon exchange approximation. Only the $\bar{3}_{c}$ representation of $SU_{c}(3)$ is considered in the present work; the $6_{c}$ representation is not considered. The tetraquarks considered are composed of a diquark and an antidiquark in color $\bar{3}$ and $3$ configuration.
Figure 1: Schematic diagram of the tetraquarks in the diquark-antidiquark picture.
In the diquark picture, the state of tetraquark is denoted as
$\left((q_{1}q^{\prime}_{1})^{{\bar{3}_{c}}}_{n_{1}^{2s_{1}+1}l_{1j_{1}}}(\bar{q}_{2}\bar{q}^{\prime}_{2})^{{{3}_{c}}}_{n_{2}^{2s_{2}+1}l_{2j_{2}}}\right)^{1_{c}}_{N^{2j+1}L_{J}},$ (1)
[…]
II.2 SSE
The spinless Salpeter equation Godfrey:1985xj ; Ferretti:2019zyh ; Bedolla:2019zwg ; Durand:1981my ; Durand:1983bg ; Lichtenberg:1982jp ; Jacobs:1986gv reads as
$M\Psi_{d,t}({{\bf r}})=\left(\omega_{1}+\omega_{2}\right)\Psi_{d,t}({{\bf r}})+V_{d,t}\Psi_{d,t}({{\bf r}}),$ (2)
where $M$ is the bound state mass (diquark or tetraquark). $\Psi_{d,t}({{\bf r}})$ are the diquark wave function and the tetraquark wave function, respectively. $V_{d,t}$ denotes the diquark potential and the tetraquark potential, respectively (see Eq. (4)). $\omega_{1}$ is the relativistic energy of constituent $1$ (quark or diquark), and $\omega_{2}$ is of constituent $2$ (quark or antidiquark),
$\omega_{i}=\sqrt{m_{i}^{2}+{\bf p}^{2}}=\sqrt{m_{i}^{2}-\Delta}\;\;(i=1,2).$ (3)
$m_{1}$ and $m_{2}$ are the effective masses of constituent $1$ and $2$, respectively.
When using diquark in multiquark systems, the interactions between quark and quark, diquark and quark, diquark and diquark are needed. In Refs. Faustov:2021hjs , these interactions are constructed with the help of the off-mass-shell scattering amplitude, which is projected onto the positive energy states. The interactions also can be established by expanding the interactions of the quark-antiquark system to the quark-quark system, and then to the diquark-antidiquark systems or the diquark-quark systems Lundhammar:2020xvw . Furthermore, the effect of the finite size of diquark is treated differently. In Refs. Faustov:2021hjs , the size of diquark is taken into account through corresponding form factors. At times, diquark is taken as being pointlike Ferretti:2019zyh ; Lundhammar:2020xvw , and we use this approximation in the present work.
Following Refs. Ferretti:2019zyh ; Bedolla:2019zwg ; Ferretti:2011zz ; Eichten:1974af , we employ the potential
$V_{d,t}$ $=-\frac{3}{4}\left[V_{c}+{\sigma}r+C\right]\left({\bf{F}_{i}}\cdot{\bf{F}_{j}}\right)_{d,t},$ (4)
where $V_{c}\propto{1/r}$ is a color Coulomb potential or a Coulomb-like potential due to one-gluon-exchange. $\sigma$ is the string tension. $C$ is a fundamental parameter Gromes:1981cb ; Lucha:1991vn . The part in the bracket is the Cornell potential Eichten:1974af . ${\bf{F}_{i}}\cdot{\bf{F}_{j}}$ is the color-Casimir,
$\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{d}\rangle=-\frac{2}{3},\quad\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{t}\rangle=-\frac{4}{3}.$ (5)
II.3 Regge trajectory relations for the heavy-heavy systems
For the heavy-heavy systems, $m_{1},m_{2}{\gg}{{|\bf p|}}$, Eq. (2) reduces to
$M\Psi_{d,t}({{\bf r}})$ $=$ $\left[(m_{1}+m_{2})+\frac{{{\bf p}}^{2}}{2\mu}\right]\Psi_{d,t}({{\bf r}})$ (6)
$+V_{d,t}\Psi_{d,t}({{\bf r}}),$
where
$\mu=m_{1}m_{2}/(m_{1}+m_{2}).$ (7)
By employing the Bohr-Sommerfeld quantization approach Brau:2000st and using Eqs. (4) and (6), the orbital and radial Regge trajectories can be obtained,
$M{\sim}\frac{3}{2}\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3}l^{2/3},\quad M{\sim}\left(\frac{3\pi}{2}\right)^{2/3}\left(\frac{\sigma^{\prime 2}}{2\mu}\right)^{1/3}{n_{r}}^{2/3},$ (8)
where
$\sigma^{\prime}=\left\{\begin{array}[]{cc}\sigma/2,&\text{diquarks},\\ \sigma,&\text{tetraquarks}.\end{array}\right.$ (9)
Consider the constant terms and the omitted terms and use Eq. (8), then we obtain Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}(x+c_{0x})^{2/3}\,\,(x=l,\,n_{r}),$ (10)
with
$\beta_{x}=c_{fx}c_{x}c_{c},\quad m_{R}=m_{1}+m_{2}+C^{\prime},$ (11)
where
$C^{\prime}=\left\{\begin{array}[]{cc}C/2,&\text{diquarks},\\ C,&\text{tetraquarks}.\end{array}\right.$ (12)
$c_{x}$ and $c_{c}$ are
$c_{c}=\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3},\quad c_{l}=\frac{3}{2},\quad c_{n_{r}}=\frac{\left(3\pi\right)^{2/3}}{2}.$ (13)
Both $c_{fl}$ and $c_{fn_{r}}$ are equal theoretically to one and are fitted in practice. In Eq. (10), $m_{1}$, $m_{2}$, $c_{x}$ and $\sigma$ are universal for the heavy-heavy systems. $c_{fx}$ and $c_{0x}$ are determined by fit.
II.4 Regge trajectory relations for the heavy-light systems
For the heavy-light systems ($m_{1}\to\infty$ and $m_{2}\to 0$), Eq. (2) simplifies to
$M\Psi_{d,t}({{\bf r}})=\left[m_{1}+{{|\bf p|}}+V_{d,t}\right]\Psi_{d,t}({{\bf r}}).$ (14)
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , we obtain from Eq. (14)
$M{\sim}2\sqrt{\sigma^{\prime}}\sqrt{l},\;M{\sim}\sqrt{2{\pi}\sigma^{\prime}}\sqrt{n_{r}},$ (15)
where $\sigma^{\prime}$ is in Eq. (9). Using Eq. (15), the parameterized formula can be written as Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}\sqrt{x+c_{0x}}\;(x=l,\,n_{r}).$ (16)
[In this subsection and the preceding subsection, the notations $m_{R}$, $\beta_{x}$, $c_{fx}$, $c_{x}$ and $c_{c}$ have the same form but different values. This should not cause confusion.] The parameter in Eq. (16) reads as
$\beta_{x}=c_{fx}c_{x}c_{c}$ (17)
with
$c_{c}=\sqrt{\sigma^{\prime}},\quad c_{l}=2,\quad c_{n_{r}}=\sqrt{2\pi}.$ (18)
Both $c_{fl}$ and $c_{fn_{r}}$ are theoretically equal to one but are fitted in practice. For the heavy-light systems, the common choice of $m_{R}$ is Selem:2006nd ; Chen:2021kfw ; Jakhad:2023mni ; Chen:2014nyo ; Chen:2017fcs ; Veseli:1996gy ; Jia:2018vwl
$m_{R}=m_{1}.$ (19)
$m_{R}$, $c_{x}$ and $\sigma$ are universal for the heavy-light systems. $c_{fx}$ and $c_{0x}$, which vary with different Regge trajectories, are determined by fitting the given Regge trajectory.
The usual Regge trajectory, Eq. (16) with (19), is obtained in the limit $m_{1}\to\infty$ and $m_{2}\to 0$. There are different ways of including the light constituent’s mass Selem:2006nd ; Nielsen:2018uyn ; Sonnenschein:2018fph ; MartinContreras:2020cyg ; Chen:2023cws ; Chen:2022flh ; Chen:2023ngj ; Chen:2014nyo ; Afonin:2014nya . Two modified formulas are proposed in Ref. Chen:2023cws , which can universally describe both the heavy-light mesons and the heavy-light diquarks. One is Eq. (16) with
$m_{R}=m_{1}+m_{2}+C^{\prime},$ (20)
where $C^{\prime}$ is in (12). Another reads
$M=m_{R}+\sqrt{\beta_{x}^{2}(x+c_{0x})+\kappa_{x}m^{3/2}_{2}(x+c_{0x})^{1/4}}$ (21)
if $m_{2}{\ll}M$, where
$m_{R}=m_{1}+C^{\prime},\quad\kappa_{x}=\frac{4}{3}\sqrt{{\pi}\beta_{x}},$ (22)
where $\beta_{x}$ is in (17). Eq. (16) with (20) is an extension of Afonin:2014nya
$M=m_{1}+m_{2}+\sqrt{a(n_{r}+{\alpha}l+b)},$ (23)
while (21) with (22) is based on the results in Selem:2006nd ; Sonnenschein:2018fph . As $m_{2}=0$, these two modified formulas, Eq. (16) with (20) and Eq. (21) with (22), become identical. As $m_{2}=0$ and $C$ is neglected, these two modified formulas reduce to the usual Regge trajectory formula for the heavy-light systems, i.e., (16) with (19). Although they give different behavior of $m_{2}$, Eq. (16) with (20) and Eq. (21) with (22) produce consistent results for $l,\,n_{r}<10$ and have the same behavior $M{\sim}x^{1/2}$ Chen:2023cws .
Table 1: The coefficients.
Heavy-heavy systems Heavy-light sytems
$\nu$ $2/3$ $1/2$
$c_{c}$ $\left({\sigma^{\prime 2}}/{\mu}\right)^{1/3}$ $\sqrt{\sigma^{\prime}}$
$c_{l}$ $3/2$ $2$
$c_{n_{r}}$ ${\left(3\pi\right)^{2/3}}/{2}$ $\sqrt{2\pi}$
In this work, Eq. (16) with (20) is employed to discuss the heavy-light systems. By summarising Eqs. (10) and (16) with (20), we have Chen:2022flh
$M=$ $m_{R}+\beta_{x}(x+c_{0x})^{\nu}\,\,(x=l,\,n_{r}),$
$m_{R}=$ $m_{1}+m_{2}+C^{\prime},\quad\beta_{x}=c_{fx}c_{x}c_{c},$ (24)
where ${\nu}$, $c_{x}$ and $c_{c}$ are listed in Table 1. $c_{fx}$ and $c_{0x}$ vary with different Regge trajectories. Eq. (II.4) can used to discuss the heavy-heavy and heavy-light systems: diquarks, mesons, baryons and tetraquarks Chen:2023web ; Chen:2023djq .
II.5 Regge trajectory relations for the hidden bottom tetraquarks
A hidden bottom tetraquark consists of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$. Both $(bq)$ and $(\bar{b}\bar{q}^{\prime})$ are heavy because the bottom quark is heavy. Therefore, the hidden bottom tetraquarks are the heavy-heavy systems in case of the $\lambda$-mode excitations between diquark and antidiquark. Meanwhile, the Regge trajectories for hidden bottom tetraquarks exhibit the properties of the heavy-light systems in case of the $\rho$-mode excitations in diquark and antidiquark. According to discussions in the preceding subsections, using Eq. (II.4) and coefficients in Table 1, we obtain the Regge trajectory relation for the hidden bottom tetraquarks
$M$ $=m_{R{\lambda}}+\beta_{x_{\lambda}}(x_{\lambda}+c_{0x_{\lambda}})^{2/3}\;(x_{\lambda}=L,\,N_{r}),$
$M_{\rho_{1}}$ $=m_{R\rho_{1}}+\beta_{x_{\rho_{1}}}\sqrt{x_{\rho_{1}}+c_{0x_{\rho_{1}}}}\;(x_{\rho_{1}}=l_{1},\,n_{r_{1}}),$
$M_{\rho_{2}}$ $=m_{R\rho_{2}}+\beta_{x_{\rho_{2}}}\sqrt{x_{\rho_{2}}+c_{0x_{\rho_{2}}}}\;(x_{\rho_{2}}=l_{2},\,n_{r_{2}}),$ (25)
where
$m_{R{\lambda}}$ $=M_{\rho_{1}}+M_{\rho_{2}}+C,$
$m_{R\rho_{1}}$ $=m_{b}+m_{q}+C/2,$
$m_{R\rho_{2}}$ $=m_{b}+m_{q^{\prime}}+C/2,$
$\beta_{L}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fL},$
$\beta_{N_{r}}$ $=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fN_{r}},$
$\mu_{\lambda}$ $=\frac{M_{\rho_{1}}M_{\rho_{2}}}{M_{\rho_{1}}+M_{\rho_{2}}},$
$\beta_{l_{1}}$ $=2\sqrt{\frac{\sigma}{2}}c_{fl_{1}},\quad\beta_{n_{r_{1}}}=\sqrt{\pi\sigma}c_{fn_{r_{1}}},$
$\beta_{l_{2}}$ $=2\sqrt{\frac{\sigma}{2}}c_{fl_{2}},\quad\beta_{n_{r_{2}}}=\sqrt{\pi\sigma}c_{fn_{r_{2}}}.$ (26)
There are three series of Regge trajectories for the hidden bottom tetraquarks, see Eq. (II.5).
III Regge trajectories for the hidden bottom tetraquarks
In this section, both the $\lambda-$trajectories and the $\rho-$trajectories for the hidden bottom tetraquarks are investigated.
III.1 Parameters
The quark masses, the string tension $\sigma$ and the parameter $C$ are from Ref. Faustov:2021hjs . The parameters for the heavy-light diquarks $(bu)$ and $(bs)$ are from Ref. Chen:2023cws and listed in Table 2. With these parameters determined, the $\rho-$modes and the diquark masses can be discussed, see Eqs. (II.5) and (II.5). To discuss the masses of the hidden bottom tetraquarks and the $\lambda-$modes excited states, the parameters $c_{fL}$, $c_{fN_{r}}$, $c_{0{L}}$ and $c_{0N_{r}}$ should be determined, see Eqs. (29) and (30). More discussions are in section C. Once all parameters are determined, the Regge trajectories can be discussed and can be used to estimate the masses of the hidden bottom tetraquarks.
Table 2: The values of parameters Chen:2023cws ; Faustov:2021hjs .
$m_{u,d}=0.33\;{{\rm GeV}}$,   $m_{s}=0.50\;{{\rm GeV}}$,   $m_{b}=4.88\;{{\rm GeV}}$,
$\sigma=0.18\;{{\rm GeV}^{2}}$,  $C=-0.3\;{{\rm GeV}}$,
$(bu)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.988$,  $c_{fl_{1}}=c_{fl_{2}}=0.965$,
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.125$,
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.18$,
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.155$,
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.22$,
$(bs)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.953$,  $c_{fl_{1}}=c_{fl_{2}}=0.919$,
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.08$,
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.115$.
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.11$,
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.16$,
III.2 $\lambda-$ and $\rho-$trajectories
Table 3: The masses of the $\lambda$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $n=u,\,d$. Eqs. (II.5), (29) and (30) are used.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ $(bn)(\bar{b}\bar{n})$ $(bn)(\bar{b}\bar{s})$ $(bs)(\bar{b}\bar{s})$
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 10.46 10.57 10.68
$|1^{1}s_{0},1^{1}s_{0},2^{1}S_{0}\rangle$ 10.91 11.02 11.12
$|1^{1}s_{0},1^{1}s_{0},3^{1}S_{0}\rangle$ 11.21 11.32 11.43
$|1^{1}s_{0},1^{1}s_{0},4^{1}S_{0}\rangle$ 11.47 11.57 11.68
$|1^{1}s_{0},1^{1}s_{0},5^{1}S_{0}\rangle$ 11.70 11.80 11.91
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 10.46 10.57 10.68
$|1^{1}s_{0},1^{1}s_{0},1^{1}P_{1}\rangle$ 10.78 10.89 11.00
$|1^{1}s_{0},1^{1}s_{0},1^{1}D_{2}\rangle$ 11.01 11.12 11.22
$|1^{1}s_{0},1^{1}s_{0},1^{1}F_{3}\rangle$ 11.20 11.31 11.41
$|1^{1}s_{0},1^{1}s_{0},1^{1}G_{4}\rangle$ 11.37 11.48 11.59
$|1^{1}s_{0},1^{1}s_{0},1^{1}H_{5}\rangle$ 11.53 11.64 11.74
Figure 2: The $\lambda-$trajectories for the hidden bottom tetraquarks. Circles represent the predicted data and the blue lines are the $\lambda-$trajectories, see Eq. (II.5). Data are listed in Table 3.
Table 4: Comparison of theoretical predictions for the masses of the hidden bottom tetraquarks (in GeV). $C$ is defined only for $q=q^{\prime}$.
$J^{PC}$ $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ Tetraquark Our FGM Faustov:2021hjs FS Ferretti:2020ewe TR Tiwari:2022azj HK Hadizadeh:2015cvx
$0^{++}$ $|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ $(bn)(\bar{b}\bar{n})$ 10.46 10.471 10.410
$(bn)(\bar{b}\bar{s})$ 10.57 10.572 10.407 10.522
$(bs)(\bar{b}\bar{s})$ 10.68 10.662 10.711 10.613
$|1^{1}s_{0},1^{1}s_{0},2^{1}S_{0}\rangle$ $(bn)(\bar{b}\bar{n})$ 10.91 10.917 10.899
$(bn)(\bar{b}\bar{s})$ 11.02 11.018 10.909 11.085
$(bs)(\bar{b}\bar{s})$ 11.12 11.111 11.244 11.100
$1^{--}$ $|1^{1}s_{0},1^{1}s_{0},1^{1}P_{1}\rangle$ $(bn)(\bar{b}\bar{n})$ 10.78 10.807 10.806
$(bn)(\bar{b}\bar{s})$ 10.89 10.907 10.790 11.024
$(bs)(\bar{b}\bar{s})$ 11.00 11.002 11.184 11.009
The tetraquarks with hidden bottom include the neutral states $(bu)(\bar{b}\bar{u})$ and $(bd)(\bar{b}\bar{d})$, the charged states $(bu)(\bar{b}\bar{d})$ and $(bd)(\bar{b}\bar{u})$, the states with open strangeness $(bs)(\bar{b}\bar{n})$ and $(bn)(\bar{b}\bar{s})$ ($n=u,\,d$), and the states with hidden strangeness $(bs)(\bar{b}\bar{s})$. We assume $m_{u}=m_{d}$, therefore, the masses for diquarks $(bd)$ and $(bu)$ are equal and it holds for the tetraquarks as well.
When calculating the $\lambda$-mode excitations, the scalar diquark and the scalar antidiquark are used. The radially and orbitally excited states of the $\lambda-$mode are calculated by using Eqs. (II.5), (29) and (30) and parameters in Table 2. The calculated results are listed in Table 3. More results can be easily obtained in a similar manner. The $\lambda-$trajectories for the hidden bottom tetraquarks are shown in Fig. 2.
Table 5: The masses of the $\rho$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $(M^{\prime},\,M^{\prime\prime})$ is for $\left(|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle,\;|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\right)$. For $(bn)(\bar{b}\bar{n})$ and $(bs)(\bar{b}\bar{s})$, $M^{\prime}=M^{\prime\prime}$. $n=u,\,d$. Eqs. (II.5), (29) and (30) are used.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ $(bn)(\bar{b}\bar{n})$ $(bn)(\bar{b}\bar{s})$ $(bs)(\bar{b}\bar{s})$
$M^{\prime}=M^{\prime\prime}$ $(M^{\prime},\,M^{\prime\prime})$ $M^{\prime}=M^{\prime\prime}$
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 10.46 (10.57, 10.57) 10.68
$|2^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 10.98 (11.08, 11.10) 11.21
$|3^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 11.27 (11.37, 11.38) 11.49
$|4^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 11.49 (11.60, 11.61) 11.71
$|5^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 11.69 (11.79, 11.79) 11.90
$|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 10.49 (10.60, 10.60) 10.71
$|2^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 10.99 (11.09, 11.11) 11.22
$|3^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 11.27 (11.38, 11.39) 11.50
$|4^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 11.50 (11.61, 11.61) 11.72
$|5^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 11.69 (11.80, 11.80) 11.91
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 10.44 (10.55, 10.55) 10.66
$|1^{1}p_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 10.82 (10.93, 10.94) 11.05
$|1^{1}d_{2},1^{1}s_{0},1^{5}S_{2}\rangle$ 11.04 (11.15, 11.16) 11.26
$|1^{1}f_{3},1^{1}s_{0},1^{7}S_{3}\rangle$ 11.22 (11.32, 11.33) 11.43
$|1^{1}g_{4},1^{1}s_{0},1^{9}S_{4}\rangle$ 11.37 (11.47, 11.47) 11.58
$|1^{1}h_{5},1^{1}s_{0},1^{11}S_{5}\rangle$ 11.50 (11.60, 11.59) 11.70
$|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 10.47 (10.58, 10.59) 10.69
$|1^{3}p_{2},1^{1}s_{0},1^{5}S_{2}\rangle$ 10.83 (10.94, 10.95) 10.06
$|1^{3}d_{3},1^{1}s_{0},1^{7}S_{3}\rangle$ 11.05 (11.16, 11.17) 11.27
$|1^{3}f_{4},1^{1}s_{0},1^{9}S_{4}\rangle$ 11.22 (11.33, 11.33) 11.44
$|1^{3}g_{5},1^{1}s_{0},1^{11}S_{5}\rangle$ 11.37 (11.48, 11.47) 11.58
$|1^{3}h_{6},1^{1}s_{0},1^{13}S_{6}\rangle$ 11.50 (11.61, 11.60) 11.71
Figure 3: The $\rho-$trajectories for the hidden bottom tetraquarks. Circles indicate the predicted data and the blue lines are the $\rho-$trajectories, see Eq. (II.5). Data are listed in Table 5.
As discussed in Section C, the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with the masses of diquark and antidiquark, $M_{\rho_{1}}$ and $M_{\rho_{2}}$, see Eqs. (II.5), (29) and (30). This means that $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with different states of diquark or antidiquark as we discuss the $\rho-$trajectories, even though $\lambda-$mode remains ground state. Using Eqs. (II.5), (29) and (30) and parameters in Table 2, the masses of the $\rho-$mode excited states are calculated and listed in Table 5.
The mass of the state $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ for $(bq)(\bar{b}\bar{q^{\prime}})$ is equal to that of the state $|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle$ for $(bq^{\prime})(\bar{b}\bar{q})$. For the $\rho-$mode excited states, the mass of the state $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ with definite parity and charge parity (for $q=q^{\prime}$) $1/\sqrt{2}\big{[}(bq)(\bar{b}\bar{q^{\prime}})\pm(bq^{\prime})(\bar{b}\bar{q})\big{]}$ is equal to that of state
$\frac{1}{\sqrt{2}}\Big{(}|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$
±$\pm|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\Big{)}$ (27)
for $(bq)(\bar{b}\bar{q^{\prime}})$ if $n_{1}^{2s_{1}+1}l_{1j_{1}}$ is not identical to $n_{2}^{2s_{2}+1}l_{2j_{2}}$. The corresponding masses should be $M=(M^{\prime}+M^{\prime\prime})/2$, where $M^{\prime}$ and $M^{\prime\prime}$ are listed in Table 5. Therefore, the corresponding tetraquarks will be degenerate in mass. The $\rho-$trajectories are shown in Fig. 3.
The obtained results for the $\lambda-$mode excited states are in agreement with other theoretical predictions, see Table 4. To our knowledge, the highly radial and orbital excited states of the $\lambda-$mode states, $|1^{1}s_{0},1^{1}s_{0},4^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},5^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}F_{3}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}G_{4}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}H_{5}\rangle$, and the $\rho-$trajectories of the hidden bottom tetraquarks are discussed for the first time.
We estimate the masses of the hidden bottom tetraquarks for the highly excited states and the masses for the more highly excited states can be calculated easily by the Regge trajectory formula. Although the highly excited states will be unstable, the calculated masses are useful. This is because various models or approaches to calculate the masses of tetraquarks often use different parameter values. These models, despite using different parameter values, often give agreeable predictions for the ground state and the lower excited states. The theoretical predictions for the highly excited states are expected to show differences, which will be instructive.
III.3 Discussions
The Regge trajectories take different form and behave differently in various energy regions Chen:2022flh ; Chen:2021kfw . The $\rho$-mode excitations are in the diquark or in the antidiquark. Both the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q})$ are the heavy-light systems; therefore, the $\rho$-trajectories of the hidden bottom tetraquarks behave as $M{\sim}x_{\rho}^{1/2}$, see Eq. (II.5). The $\lambda$-mode excitations are between diquark $(bq)$ and antidiquark $(\bar{b}\bar{q})$, with the tetraquark $(bq)(\bar{b}\bar{q})$ being the heavy-heavy system; hence, the $\lambda$-trajectories of the hidden bottom tetraquarks behave as $M{\sim}x_{\lambda}^{2/3}$. Discussions on other types of tetraquarks can be found in Ref. plan .
According to Eq. (II.5), there are three series of Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. There are degeneracy in the $\rho$-trajectories, corresponding to the masses of the $\rho$-mode excited states, as discussed in subsection III.2.
The curvature of the Regge trajectories holds significant importance Chen:2018bbr . In potential models, the curvature is related to the dynamic equation and the confining potential. In Ref. Chen:2023web , we suggest that all Regge trajectories for the diquarks, mesons, baryons and tetraquarks are concave downwards in the $(M^{2},\,x)$ $(x=L,\,N_{r})$ planes. For the light systems, the Regge trajectories are also concave when considering the masses of the light constituents. In fact, the trajectories of baryons and tetraquarks discussed in Refs. Chen:2023web ; Chen:2023djq are the $\lambda$-trajectories. In this work, we show that for the hidden bottom tetraquarks, not only the $\lambda$-trajectories but also the $\rho$-trajectories are concave in the $(M^{2},\,x)$ planes.
IV Conclusions
In this work, we obtain the Regge trajectory formulas for the hidden bottom tetraquarks by using the Regge trajectory formula for the heavy-heavy systems and the newly proposed Regge trajectory relation for the heavy-light diquarks. We apply the Regge trajectory formulas to investigate the hidden bottom tetraquarks which are composed of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks with the hidden bottom are estimated, and the $\lambda$-trajectories and $\rho$-trajectories are discussed.
For the hidden bottom tetraquarks, there are three series of the Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. They exhibit different behaviors. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ because the $\rho$-mode excitations are in diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$, which are the heavy-light systems. In contrast, the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ because the $\lambda$-mode excitations occur between the diquark and the antidiquark, with the the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ forming a heavy-heavy system $(bq)(\bar{b}\bar{q}^{\prime})$. Moreover, not only the $\lambda$-trajectories but also the $\rho$-trajectories of the hidden bottom tetraquarks are concave downwards in the $(M^{2},\,x)$ plane.
Appendix A States of diquarks
Table 6: The completely antisymmetric states for the diquarks in $\bar{3}_{c}$ Feng:2023txx . $j$ is the spin of the diquark $(qq^{\prime})$, $s$ denotes the total spin of two quarks, $l$ represents the orbital angular momentum. $n=n_{r}+1$, $n_{r}$ is the radial quantum number, $n_{r}=0,1,2,\cdots$.
Spin of diquark Parity Wave state Configuration
( $j$ ) $(j^{P})$ $(n^{2s+1}l_{j})$
j=0 $0^{+}$ $n^{1}s_{0}$ $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{1}s_{0}}$
$0^{-}$ $n^{3}p_{0}$ $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{3}p_{0}}$
j=1 $1^{+}$ $n^{3}s_{1}$, $n^{3}d_{1}$$\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n^{3}s_{1}}$,  $\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n^{3}d_{1}}$
$1^{-}$ $n^{1}p_{1}$, $n^{3}p_{1}$$\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n^{1}p_{1}}$,  $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{3}p_{1}}$
j=2 $2^{+}$ $n^{1}d_{2}$, $n^{3}d_{2}$ $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{1}d_{2}}$,  $\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n^{3}d_{2}}$
$2^{-}$ $n^{3}p_{2}$, $n^{3}f_{2}$ $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{3}p_{2}}$,  $[qq^{\prime}]^{{\bar{3}_{c}}}_{n^{3}f_{2}}$
$\cdots$ $\cdots$ $\cdots$ $\cdots$
In this section, we list the completely antisymmetric states for the diquarks in $\bar{3}_{c}$ Feng:2023txx .
Appendix B States of tetraquarks
The states of tetraquark in the diquark picture are listed in Table 7.
Table 7: The states of tetraquark in the diquark picture. The diquark notation is in 6. Here, $q$ and $q^{\prime}$ represent both the light quarks and the heavy quarks. The superscript $1_{c}$ is omitted.
$J^{P}$ $(l_{1},l_{2},L)$ Configuration
$0^{+}$ $(0,0,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{1}s_{0}}\right)_{N^{1}S_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{1}S_{0}}$,
$(1,1,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{0}}\right)_{N^{1}S_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{1}S_{0}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{1}S_{0}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{2}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{1}S_{0}}$,
($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{1}S_{0}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{1}S_{0}}$
$(0,1,1)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}P_{0}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}P_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{0}}\right)_{N^{3}P_{0}}$,
($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}P_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}P_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{3}P_{0}}$
$(1,0,1)$ ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{0}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{0}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{0}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$,
($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{2}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$
$\cdots$ $\cdots$
$0^{-}$ $(0,0,1)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{0}}$,
$(1,0,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{1}s_{0}}\right)_{N^{1}S_{0}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{1}S_{0}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{1}S_{0}}$,
$(0,1,0)$ $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{0}}\right)_{N^{1}S_{0}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{1}p_{1}}\right)_{N^{1}S_{0}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{1}}\right)_{N^{1}S_{0}}$,
$\cdots$ $\cdots$
$1^{+}$ $(0,0,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{1}s_{0}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$,
$(1,1,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,
($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{0}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{3}S_{1}}$,
$\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{0}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{3}S_{1}}$,
$\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{2}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{2}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{2}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{3}S_{1}}$,
$(0,1,1)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}P_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{0}}\right)_{N^{3}P_{1}}$,
($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{3}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{3}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{3}P_{1}}$
($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{1}p_{1}}\right)_{N^{5}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{1}}\right)_{N^{5}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{3}p_{2}}\right)_{N^{5}P_{1}}$
$(1,0,1)$ ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{0}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$,
($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{2}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$
($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{5}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{5}P_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{3}p_{2}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{3}s_{1}}\right)_{N^{5}P_{1}}$
$\cdots$ $\cdots$
$1^{-}$ $(0,0,1)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{3}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}P_{1}}$,
$\left([qq^{\prime}]^{3_{c}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{1}s_{0}}\right)_{N^{1}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{1}P_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{5}P_{1}}$,
$(1,0,0)$ $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{0}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{{\bar{3}_{c}}}_{n_{1}^{1}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{1}s_{0}}\right)_{N^{3}S_{1}}$,
$\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{1}}[\bar{q}\bar{q}^{\prime}]^{3_{c}}_{n_{2}^{1}s_{0}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{{\bar{3}_{c}}}_{n_{1}^{3}p_{2}}\{\bar{q}\bar{q}^{\prime}\}^{3_{c}}_{n_{2}^{3}s_{1}}\right)_{N^{3}S_{1}}$
$(0,1,0)$ ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{0}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{1}s_{0}}\{\bar{q}\bar{q}^{\prime}\}^{{\bar{3}_{c}}}_{n_{2}^{1}p_{1}}\right)_{N^{3}S_{1}}$,
($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,  $\left([qq^{\prime}]^{3_{c}}_{n_{1}^{1}s_{0}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{1}}\right)_{N^{3}S_{1}}$,  ($\left(\{qq^{\prime}\}^{3_{c}}_{n_{1}^{3}s_{1}}[\bar{q}\bar{q}^{\prime}]^{{\bar{3}_{c}}}_{n_{2}^{3}p_{2}}\right)_{N^{3}S_{1}}$
$\cdots$ $\cdots$
Appendix C Determination of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes
In this section, we determine the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes of the hidden bottom tetraquarks. Apply Eq. (II.4) to fit the Regge trajectories for the doubly heavy mesons, the $\lambda$-modes of heavy-heavy baryons composed of one heavy quark and one doubly heavy (or one heavy-light) diquark and the $\lambda$-modes of heavy-heavy tetraquarks composed of one doubly heavy (or heavy-light) diquark and one doubly heavy (or heavy-light) antidiquark. The quality of a fit is measured by the quantity $\chi^{2}$ defined by Sonnenschein:2014jwa
$\chi^{2}=\frac{1}{N-1}\sum^{N}_{i=1}\left(\frac{M_{fi}-M_{ei}}{M_{ei}}\right)^{2},$ (28)
where $N$ is the number of points on the trajectory, $M_{fi}$ is fitted value and $M_{ei}$ is the experimental value or the theoretical value of the $i$-th particle mass. The parameters are determined by minimizing $\chi^{2}$.
Table 8: The spin averaged masses of the radially excited states of mesons and baryons, and the masses of tetraquarks (in ${{\rm GeV}}$).
$1S$ $2S$ $3S$ $4S$ $5S$ $6S$
$c\bar{c}$ $3.0687$ $3.6740$ $4.0273$ $4.4115$ $4.8305$ $5.1640$
$c\bar{b}$ 6.3184 6.8793 7.2500 7.6030 7.9430
$b\bar{b}$ 9.4450 10.0173 10.3486 10.5778 10.8645 11.0812
$\Omega_{ccb}$ 7.9940 8.4097
$\Omega_{cbb}$ 11.2107 11.6967
$[cs][\bar{c}\bar{s}]$ 4.051 4.604
$[bs][\bar{b}\bar{s}]$ 10.662 11.111
Table 9: The spin averaged masses of the orbitally excited states of mesons and baryons, and the masses of tetraquarks (in ${{\rm GeV}}$).
$1S$ $1P$ $1D$ $1F$ $1G$ $1H$
$c\bar{c}$ 3.0687 3.5253 3.8186 4.0720 4.3435 4.5931
$c\bar{b}$ 6.3184 6.7486 7.0261 7.2720 7.4879
$b\bar{b}$ 9.4450 9.8997 10.1629 10.3467 10.5127 10.6711
$\Omega_{ccb}$ 7.9940 8.2643 8.4741
$\Omega_{cbb}$ 11.2107 11.5261 11.8053
$[cs][\bar{c}\bar{s}]$ 4.051 4.466 4.728
$[bs][\bar{b}\bar{s}]$ 10.662 11.002 11.216
The masses used are listed in Tables 8 and 9. For the mesons and baryons, the listed data are the spin averaged masses. If the corresponding states are experimentally determined, the PDG data from Ref. pdg:2024 are used. For the undetermined states, the theoretical data from Ref. Ebert:2011jc ; Faustov:2021qqf are used. (In Ref. Ebert:2011jc , the mass of the $\Omega_{cbb}$ $(7/2)^{+}$(1D) state is not provided, so we use the value $11.807$ ${{\rm GeV}}$.) The theoretical data for the tetraquarks are from Faustov:2021hjs ; Faustov:2022mvs . Beside the bound states’ masses, some parameters are listed in Table 2. The masses of the axial vector diquarks are $m_{\{cc\}}=3.14$ ${{\rm GeV}}$, $m_{\{bb\}}=9.63$ ${{\rm GeV}}$ Feng:2023txx . The masses of the scalar diquarks $[cs]$ and $[bs]$ are calculated by using the parameters in Table 2 and Eq. (II.4) or (II.5) Chen:2023cws .
The fitted parameters are listed in Table 10. The fitted Regge trajectories are plotted in the $((M-m_{R})^{3/2},\;x)$ plane, see Fig. 4. The results are consistent with those in Ref. Chen:2023djq .
The fitted $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary across different Regge trajectories for different states and for different systems, see Fig. 5 and Table 10. In case of the baryons and tetraquarks, these parameters are sensitive to diquark masses. For example, $c_{fN_{r}}=1.218$ when $[bc]=6.38$ ${{\rm GeV}}$ Feng:2023txx while $c_{fN_{r}}=1.027$ when $[bc]=6.519$ ${{\rm GeV}}$, as determined by fitting the radial Regge trajectory for the tetraquark $[bc][\bar{b}\bar{c}]$ Faustov:2022mvs .
According to Eq. (II.5), $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ are needed to determine a Regge trajectory as $m_{R_{\lambda}}$ can be calculated by using Eq. (II.5) and parameters in Table 2. Two or more states on the Regge trajectory are needed to obtain $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$. Since the hidden bottom tetraquarks are not found experimentally, we choose $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ by fitting these parameters from other systems. By fitting the data in Table 10, we have
$c_{fL}=$ $1.116+0.013\mu_{\lambda},\;c_{0L}=0.540-0.141\mu_{\lambda},$ (29)
$c_{fN_{r}}=$ $1.008+0.008\mu_{\lambda},\;c_{0N_{r}}=0.334-0.087\mu_{\lambda},$ (30)
where $\mu_{\lambda}$ is the reduced masses, see Eq. (7) or (II.5). As more experimental data or more theoretical data become available, the fitted formulas will be refined.
Table 10: The fitted $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$.
$c\bar{c}$ $c\bar{b}$ $b\bar{b}$ $\Omega_{ccb}$ $\Omega_{cbb}$ $[cs][\bar{c}\bar{s}]$ $[bs][\bar{b}\bar{s}]$
$(c_{fN_{r}},\,c_{0N_{r}})$ (0.994, 0.191) (1.028, 0.135) (1.046, 0.0) (0.994, 0.333) (1.037, 0.347) (1.025, 0.255) (1.022, 0.087)
$(c_{fL},\,c_{0L})$ (1.120, 0.320) (1.154, 0.231) (1.187, 0.0) (1.047, 0.559) (1.137, 0.537) (1.151, 0.399) (1.164, 0.133)
Figure 4: The radial and orbital Regge trajectories (the black lines) for mesons, baryons and tetraquarks. The dots are the spin averaged masses of mesons and baryons and the masses of tetraquarks. The used data are listed in Tables 8, 9 and 10.
Figure 5: The fitted $c_{fx_{\lambda}}$ (fig. 5) and $c_{0x_{\lambda}}$ (fig. 5), ($x_{\lambda}=L,\,N_{r}$). The data are listed in Table 10. $\mu_{\lambda}$ is the reduced masses, see Eq. (7) or (II.5). The black filled circles (mesons), the red filled squares (tetraquarks), and the green filled triangles (baryons) are the fitted $c_{fL}$ or $c_{0L}$. The black empty circles (mesons), the red empty squares (tetraquarks), and the green empty triangles (baryons) are the fitted $c_{fN_{r}}$ or $c_{0N_{r}}$. The black lines ($c_{fL}$ and $c_{0L}$) and the dashed lines ($c_{fN_{r}}$ and $c_{0N_{r}}$) are the linear fit, see Eqs. (29) and (30).